the sum of two numbers is 53. if twice the first number minus three times the second number is 26, find both numbers.
I've tried to solve this but it doesn't look right. I've gotten x= 10.28 but I dont think that that is correct and I have yet to solve for y but I would like to make sure that the x value I got is correct or not

Answer by macston(5194)

X+Y=53; 2X-3Y=26
Solve first for X: X=53-Y
Substitute in second
2(53-Y)-3Y=26
106-2Y-3Y=26 subtract 196 from each side
-5Y=-80
Y=16
X+Y=53
X+16=53
X=37
ANSWERS the first number is 37, the second is 16